what is the answer to -4p-3=2+p? I get negatives mixed up
Answer by fractalier(6550)

From -4p - 3 = 2 + p
let us add 4p to both sides and get
-3 = 2 + 5p
now subtract two from both sides and get
-5 = 5p
now divide by five
p = -1